OpenStudy (anonymous):

let f(x)=-2x<0 0 x>0 which of the following us equivalent A.f(x)=x - absolute value x B.f(x)=2 absolute value x C.f(x)=2x D.f(x)=absolute value x-x E.f(x)=x+absolute value x

OpenStudy (anonymous):

@sidsiddhartha can u help me

OpenStudy (sidsiddhartha):

do u know what is meaning of absolute value of a function

OpenStudy (anonymous):

yea

OpenStudy (sidsiddhartha):

ok so first let's take a example \[f(x)=absolute(x)\] so \[f(x)=x,x \ge 0\]\[f(x)=-x,x \le 0\] ok ??

OpenStudy (anonymous):

yea

OpenStudy (sidsiddhartha):

so according to this theorem try option D \[f(x)=absolute(x)-x\] so \[f(x)=-x-x=-2x,x<0\] and \[f(x)=x-x=0,x>0\] so option D satisfies the conditions

OpenStudy (sidsiddhartha):

do u get it?

OpenStudy (anonymous):

yea can u hlp me with a few more

OpenStudy (anonymous):

which interval does the inequality absolute value <3 define A.(-8,3) B.(-3,3) C.(0,3) D.(3,8)

OpenStudy (sidsiddhartha):

absolute value what?

OpenStudy (anonymous):

of X sorry